An electron and a proton are each placed at rest in a uniform electric field of magnitude 528 N/C. Calculate the speed of each particle 45.2 ns after being released electron proton m/s m/s
Answer #1

F = qE, will be same on both as both have same charge
=>F = 1.6 x 10^-19 x 528 = 844.80 x 10^-19 N
=>For electron:- a(e) = F/m(e) = (844.80 x 10^-19)/(9.109×10^−31) = 9.3 x 10^13 m/s^2
=>For proton:- a(p) = F/m(p) = (844.80 x 10^-19)/(1.67×10^−27) = 5.06 x 10^10 m/s^2
For electron:-
=>By v = u + at
=>v = 0 + 9.3 x 10^13 x 45.2 x 10^-9 = 4.20 x 10^6 m/s
For proton:-
=>By v = u + at
=>v = 0 + 5.06 x 10^10 x 45.2 x 10^-9 = 2.29 x 10^3 m/s